|
|
LCLS SC Downtime JOBS 16-DEC-2025 06:00 to 21-DEC-2026 16:00 |
|
|
|
|
| 250 | 162 | 78 | 8 | 2 | 20 | 22 | 102 | 07-FEB-2026 16:47:41 | 230 |
|
|
|
INJECTOR WEST |
| Shop |
| 180438 | 15 | PEM | Active | Yes | No | DC MAGNETS | 05-JAN-2026 00:00 | 30-JUN-2026 00:00 | Lam, Briant K. | disconnect cables for bucking coil, solenoids all 5 correctors and reconnect later after the work is done | Zhou, Feng |
| 180438 | 11 | CTL | Active | Yes | No | OTHER | 05-JAN-2026 00:00 | - | Bong, James | remove tuners and reinstall them later | Zhou, Feng |
| 180438 | 12 | AEG | Active | Yes | No | OTHER | 05-JAN-2026 00:00 | - | Gassner, Georg L. | metrology correctors & SOL1/2 before the work starts and after the work is complete
metrology M1 again after | Zhou, Feng |
| 180438 | 13 | CTL | Active | Yes | No | OTHER | 05-JAN-2026 00:00 | - | Alverson, Shawn Christopher | reconnect laser control cables | Zhou, Feng |
| 180438 | 19 | CTL | Active | Yes | No | OTHER | 06-JAN-2026 00:00 | 09-JAN-2026 00:00 | Hoobler, Sonya | Reconnect cables for BPM1B and BPM2B | Zhou, Feng |
| 180438 | 1 | OTHER | Active | Yes | No | OTHER | 16-DEC-2025 00:00 | - | Zhou, Feng | Gun+LEB repair | Zhou, Feng |
| 180438 | 21 | MFDM | Active | Yes | No | OTHER | 21-JAN-2026 00:00 | 22-JAN-2026 00:00 | Correa, Daniel | Remove CM01 endcap support structure (Correa)
Open the CM01 big bellow and close the gate valve in CM01 (Coy) | Zhou, Feng |
| 180438 | 26 | MFDM | Active | Yes | No | OTHER | 01-APR-2026 00:00 | 01-APR-2026 00:00 | Correa, Daniel | Recovery for CM01 endcap structure & its accessories
, reinstall CM01 endcap support structure (MFD-Correa) | Zhou, Feng |
| 180438 | 3 | MFDM | Active | Yes | No | OTHER | 05-JAN-2026 00:00 | - | Correa, Daniel | remove gun & buncher water hoses and blow out, and reinstall later
remove the gun SS water tubing and reinsta | Zhou, Feng |
| 180438 | 6 | OTHER | Active | Yes | No | OTHER | 16-DEC-2025 00:00 | - | Gassner, Georg L. | metrology laser table, laser Mirror M1 | Zhou, Feng |
| 180438 | 7 | OTHER | Active | Yes | No | OTHER | 05-JAN-2026 00:00 | - | Benwell, Andy | disconnect gun & buncher RF probes, and reinstall later
Remove buncher WGs, and reinstall later
Reinstall g | Zhou, Feng |
| 180438 | 8 | OTHER | Active | Yes | No | OTHER | 02-FEB-2026 00:00 | 02-MAR-2026 00:00 | Reed, Dentell V. | remove loadlock transverse arm (from central chamber flange to aisle)
remove the bottom chamber of the centra | Zhou, Feng |
| 180438 | 9 | OTHER | Active | Yes | No | OTHER | 08-JAN-2026 00:00 | - | Karimian, Sherwin | disconnect NEGs and Ion pumps cables and secure them, and reconnect them later
replace cables for some hot fi | Zhou, Feng |
| 178255 | 2 | CTL | Active | Yes | No | PROF | 06-APR-2026 00:00 | 16-JAN-2027 00:00 | Curtis, Courtney A. | Replace YAG01B | Curtis, Courtney A. |
| 182230 | 1 | CTL | Active | Yes | No | PROF | 15-APR-2026 00:00 | 17-APR-2026 00:00 | Curtis, Courtney A. | Re-install YAG01b camera | Curtis, Courtney A. |
| 180438 | 27 | MFDM | Active | Yes | No | RF | 24-APR-2026 00:00 | 24-APR-2026 00:00 | Correa, Daniel | Reinstall WG support structure (MFD - Correa) | Zhou, Feng |
| 182479 | 1 | ADSO | Active | Yes | No | SAFETY | 09-FEB-2026 00:00 | 09-FEB-2026 00:00 | Van Hoover, Zoe | ADSO inspection of S0 manway alcove fall protection SRL | Van Hoover, Zoe |
| 180438 | 29 | CTL | Active | Yes | No | TOROIDS | 06-JAN-2026 00:00 | - | Murphy, James Matthew | reconnect cables for Toroid IM01 | Zhou, Feng |
| 180438 | 22 | MFDV | Active | Yes | No | VACUUM | 30-JAN-2026 00:00 | 13-FEB-2026 00:00 | Reed, Dentell V. | Replace gun window, NEGs/IPs, & RGA | Zhou, Feng |
| 180438 | 23 | MFDV | Active | Yes | No | VACUUM | 18-FEB-2026 00:00 | 25-FEB-2026 00:00 | Reed, Dentell V. | LEB works -
1. Replace the gate valves, VV01 and VV02
2. Replace 2x NEG pumps (Z400)
3. Add 1x NEG/Ion pump | Zhou, Feng |
| 180438 | 24 | MFDV | Active | Yes | No | VACUUM | 27-FEB-2026 00:00 | 18-MAR-2026 00:00 | Reed, Dentell V. | Gun & LEB bake-out -
1. Bakeout preparation (VAC/Bostic/TC/Feng)
2. cool down & Remove foils and equipment ( | Zhou, Feng |
| 180438 | 28 | CTL | Active | Yes | No | VACUUM | 03-MAR-2026 00:00 | - | Karimian, Sherwin | Re certify Vacuum PLC GUNB | Zhou, Feng |
| 181519 | 2 | CTL | Active | Yes | No | VACUUM | 13-JAN-2026 00:00 | - | Lai, Chinh Duc | Install new secondary PLC crate, power supply panel, and wire cross-connects in rack to accommodate new Gamma | Karimian, Sherwin |
| 180438 | 20 | MFDM | Active | Yes | No | WATER | 13-JAN-2026 00:00 | 16-JAN-2026 00:00 | Correa, Daniel | Remove gun water hoses & blowout (MFD - Correa)
Remove the shielding to protect gun windows (MFD - Correa)
R | Zhou, Feng |
| 180438 | 25 | MFDM | Active | Yes | No | WATER | 23-MAR-2026 00:00 | 26-MAR-2026 00:00 | Correa, Daniel | Recovery Gun + LEB -
1. reconnect SS water tubes (MFD-Correa)
2. reconnect water hoses (MFD-Correa)
3. re | Zhou, Feng |
| 180438 | 5 | CTL | Complete | Yes | No | LASER | 16-DEC-2025 00:00 | - | Alverson, Shawn Christopher | disconnect s0 laser control cable | Zhou, Feng |
| 181521 | 1 | CTL | Complete | Yes | No | MOVERS | 06-JAN-2026 00:00 | - | Bong, James | Disconnect, tag, and coil solenoids motion control cables in preparation for LDT work | Montironi, Alex |
| 181521 | 2 | CTL | Complete | Yes | No | MOVERS | 06-JAN-2026 00:00 | - | Bong, James | Disconnect, tag, and coil collimator/YAG motion cables in preparation for LDT gun rework | Montironi, Alex |
| 180438 | 10 | CBL | Complete | Yes | No | OTHER | 05-JAN-2026 00:00 | - | Boucher, Salime | disconnect cables around CM01 endcap area and organize these cables
remove the vertical cable tray
install | Zhou, Feng |
| 180438 | 14 | CTL | Complete | Yes | No | OTHER | 06-JAN-2026 00:00 | - | Hoobler, Sonya | Disconnect cables for BPM1B and BPM2B | Zhou, Feng |
| 180438 | 18 | CTL | Complete | Yes | No | OTHER | 06-JAN-2026 00:00 | - | Murphy, James Matthew | Disconnect cables for toroid IM01 | Zhou, Feng |
| 180438 | 2 | MFDM | Complete | Yes | No | OTHER | 17-DEC-2025 00:00 | - | Correa, Daniel | remove s0 laser table
| Zhou, Feng |
| 180438 | 4 | CTL | Complete | No | No | OTHER | 17-DEC-2025 00:00 | - | Lai, Chinh Duc | remove s0 gun u-shape waveguide relevant RTDs | Zhou, Feng |
| 180634 | 1 | AEG | Complete | Yes | No | OTHER | 19-DEC-2025 00:00 | - | Gassner, Georg L. | Map position of parts for the LCLS-II load lock | Gassner, Georg L. |
| 180740 | 1 | CEFAS | Complete | Yes | No | OTHER | 17-DEC-2025 00:00 | 18-DEC-2025 00:00 | Poitier, Sebastian | ASG to supply and install new temperature and humidity sensors on the outside of the laser enclosure. The sens | Poitier, Sebastian |
| 181560 | 1 | CTL | Complete | Yes | No | PROF | 06-JAN-2026 00:00 | - | Curtis, Courtney A. | Disconnect, remove/move YAG01b cables and camera system
Camera was decided to be split off as another part ne | Curtis, Courtney A. |
| 181560 | 2 | CTL | Complete | No | No | PROF | 22-JAN-2026 00:00 | - | Curtis, Courtney A. | Remove YAG01b camera system | Curtis, Courtney A. |
| 181519 | 1 | CTL | Complete | Yes | No | VACUUM | 12-JAN-2026 00:00 | - | Karimian, Sherwin | Replace SAES SIP Power controllers in L2KA00-02 and L2KA00-03 (11 total) with Gamma controllers (3 total) in t | Karimian, Sherwin |
| 182369 | 1 | CTL | Complete | Yes | No | VACUUM | 01-JAN-2026 00:00 | 30-JAN-2026 00:00 | Sanchez, Daniel | Disconnecting cables L2A02326 from the controller side in the gallery Rack L2KG00-15 to support ongoing sector | Sanchez, Daniel |
| 180229 | 1 | CTL | Drop | No | No | PROF | 09-JAN-2026 00:00 | - | Curtis, Courtney A. | Remove unused YAG01B pneumatic control cable.
Cable shop is removing cable. | Curtis, Courtney A. |
LINAC WEST (S1-9) |
| Shop |
| 182180 | 1 | SWE | Active | No | No | ARCHIVER | 30-SEP-2026 00:00 | - | Summers, Tasha | Reminder to resume PVs by end of LDT | Summers, Tasha |
| 181384 | 1 | CTL | Active | Yes | No | ATCA | 16-DEC-2025 00:00 | 09-JAN-2026 00:00 | Vu, Minh K | L2KG03-6033 P30 shm-l2b-sp02-1/5 (M300). Reset "force close" handle on slot 2 carrier. | Vu, Thuy Mong |
| 181443 | 2 | CTL | Active | Yes | No | ATCA | 19-JAN-2026 00:00 | 27-FEB-2026 00:00 | Vu, Minh K | SEC00-SEC05 Install NIC card 10G for all 8100s that connect to ATCA - SC ONLY
1. L2KA00-0542 Alcove cpu-gu | Vu, Thuy Mong |
| 178234 | 1 | SSED | Active | No | Yes | BCS | 19-JUN-2025 10:00 | 21-JUL-2025 18:00 | Cobau, Bill | Download new firmware to all of the LCLS II BCS LBLM chassis that will filter out the noise from the SC BEAM o | Cobau, Bill |
| 161602 | 1 | CTL | Active | Yes | No | DUMPS/COLLIMATORS | 20-JAN-2026 00:00 | - | Bong, James | inspect and remount collimator termination panel for CEHTR | Balakrishnan, Namrata |
| 174788 | 1 | CEFOI | Active | Yes | No | ODM | 12-JAN-2026 00:00 | 06-FEB-2026 00:00 | Delgado, Sal | Installation of 1/8" air collecting tubes for HE ODM at penetrations S08-09, S09-09 and S10-04. Tubes drop dow | Nicolas, Ludovic Y. |
| 174788 | 2 | CEFOI | Active | Yes | No | ODM | 21-JUL-2025 00:00 | 15-SEP-2026 00:00 | Delgado, Sal | LI08 move ODM controller from 08-06 to penetration 08-09 (Ch3). This includes the Oxigraf air tubing going dow | Nicolas, Ludovic Y. |
| 174788 | 3 | CEFOI | Active | Yes | No | ODM | 20-JAN-2026 00:00 | 01-JUN-2026 00:00 | Delgado, Sal | Remove Oxigraf air lines from Pen 07-11 and install replacement ones at Pen 07-10.
W.O for CEFOI - FWKD0117 | Nicolas, Ludovic Y. |
| 180520 | 1 | SSED | Active | Yes | No | ODM | 05-JAN-2026 00:00 | 01-OCT-2026 00:00 | Nicolas, Ludovic Y. | Installation of 4x4 supports in Sector 08 and Sector 09 AORs as support for ODH blue strobes. Supports tie in | Nicolas, Ludovic Y. |
| 180520 | 2 | SSED | Active | Yes | No | ODM | 05-JAN-2026 00:00 | 31-MAR-2026 00:00 | Nicolas, Ludovic Y. | Pull wires in fire conduits of Linac West Sectors 08, 09 and 10, install the strobes/horns and connect them. | Nicolas, Ludovic Y. |
| 180520 | 3 | SSED | Active | Yes | No | ODM | 05-JAN-2026 00:00 | 01-OCT-2026 00:00 | Nicolas, Ludovic Y. | Installation of 4x4 steel box support for ODH blue horn/strobe at the Sector 09 PPS entrance in Klystron Galle | Nicolas, Ludovic Y. |
| 180520 | 4 | SSED | Active | Yes | No | ODM | 05-JAN-2026 00:00 | 01-OCT-2026 00:00 | Nicolas, Ludovic Y. | Installation of one 4x4 steel box support for ODH blue strobe/horn at the PPS entrance of the S10 adit (outsid | Nicolas, Ludovic Y. |
| 180841 | 12 | SSED | Active | Yes | No | ODM | 01-OCT-2026 00:00 | 18-DEC-2026 00:00 | Nicolas, Ludovic Y. | Turn the Linac ODM System back ON on all sectors 00 to 10. | Perez, Jase |
| 180841 | 13 | SSED | Active | Yes | No | ODM | 02-NOV-2026 00:00 | 18-DEC-2026 00:00 | Nicolas, Ludovic Y. | Linac ODM System certification (IAT). | Perez, Jase |
| 180841 | 19 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install new S07 UltraFlow Duo Oxigraf, its cabinet and wire it for operations at Penetration P07-10.
The S07 | Perez, Jase |
| 180841 | 20 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install new S08 UltraFlow Duo Oxigraf, its cabinet and wire it for operations at Penetration P08-09. | Perez, Jase |
| 180841 | 21 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install new S09 UltraFlow Duo Oxigraf, its cabinet and wire it for operations at Penetration P09-09. | Perez, Jase |
| 180841 | 22 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install new S10 UltraFlow Duo Oxigraf, its cabinet and wire it for operations at Penetration P10-04. | Perez, Jase |
| 180841 | 23 | SSED | Active | Yes | No | ODM | 16-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install ODM system panels and wiring in L2KA08-08 rack | Perez, Jase |
| 180841 | 24 | SSED | Active | Yes | No | ODM | 16-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install ODM system panels and wiring in L2KA09-14 rack | Perez, Jase |
| 180841 | 25 | SSED | Active | Yes | No | ODM | 16-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install ODM system panels and wiring in L2KA10-17 rack | Perez, Jase |
| 181557 | 10 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install S05 UltraFlow Duo Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 11 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove S06 Original Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 12 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install S06 UltraFlow Duo Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 1 | SSED | Active | Yes | No | ODM | 05-JAN-2026 00:00 | 16-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove S00 original Oxigraf to allow for its swap with UltraFlow model | Nicolas, Ludovic Y. |
| 181557 | 6 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install S03 UltraFlow Duo Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 7 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove S04 Original Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 8 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install S04 UltraFlow Duo Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 9 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove S05 Original Oxigraf | Nicolas, Ludovic Y. |
| 175702 | 3 | RPFO | Active | No | No | OTHER | 09-FEB-2026 00:00 | - | Frosio, Thomas | RP inspections -
-Injector West entry maze
-S0 to S10 klystron gallery
-Inside S10 ADIT
-S10 wall SC side | Frosio, Thomas |
| 179583 | 5 | AEG | Active | No | No | OTHER | 10-FEB-2026 00:00 | - | Gassner, Georg L. | 3D scan of S7 - S10 while it is empty | Gassner, Georg L. |
| 180219 | 4 | MFDM | Active | Yes | Yes | OTHER | 05-JAN-2026 00:00 | 30-SEP-2026 00:00 | Atkinson, Richard D. | Remove BRB1 dipole and CDOG collimator and relocate it downstream by S10 wall in linac-west
Remove BRB2 dip | White, Dominique A |
| 180219 | 7 | MFDV | Active | Yes | No | OTHER | 18-DEC-2025 00:00 | 30-SEP-2026 00:00 | Diaz, Gregory L. | Disconnect and remove UHV beam-line vacuum components of the dogleg in linac west S7 through S10 wall.
Follo | White, Dominique A |
| 180459 | 1 | OTHER | Active | Yes | Yes | OTHER | 22-JAN-2026 00:00 | 13-MAR-2026 00:00 | White, Dominique A | Install 7 sets of Cm stands in housing between S7 and S8.
SOW steps -
1. Drill and install anchors (contract | Checchin, Mattia |
| 180459 | 2 | OTHER | Active | Yes | Yes | OTHER | 22-JAN-2026 00:00 | 30-APR-2026 00:00 | White, Dominique A | Install L3-L4 transition and scavenging pump anchors in S7.
SOW steps -
1. drill holes and install anchors ( | Checchin, Mattia |
| 180459 | 3 | OTHER | Active | Yes | No | OTHER | 19-JAN-2026 00:00 | 30-APR-2026 00:00 | Checchin, Mattia | Demo 4 concrete pads in the S7 housing. | Checchin, Mattia |
| 180841 | 29 | AMG | Active | Yes | Yes | OTHER | 02-FEB-2026 00:00 | 20-MAR-2026 00:00 | Perez, Jase | Project to install new duct work, support stands, etc. for new S10 adit AHU | Perez, Jase |
| 180841 | 6 | OTHER | Active | Yes | No | OTHER | 02-FEB-2026 00:00 | 27-FEB-2026 00:00 | Perez, Jase | The existing Sector 07/08 PPS Gate will get demo'd during the long downtime. | Perez, Jase |
| 181038 | 1 | MFDM | Active | Yes | No | OTHER | 02-FEB-2026 00:00 | 30-JUN-2026 00:00 | Checchin, Mattia | Install ceiling waveguides in penetrations P07-09 through P10-04 (P09-09 and P09-10 excluded) | Checchin, Mattia |
| 181038 | 2 | MFDM | Active | Yes | No | OTHER | 01-MAR-2026 00:00 | 30-SEP-2026 00:00 | Checchin, Mattia | Install vertical waveguides in penetrations P07-09 through P10-04 | Checchin, Mattia |
| 181088 | 1 | MFDM | Active | Yes | No | OTHER | 09-MAR-2026 00:00 | 28-AUG-2026 00:00 | Atkinson, Richard D. | Unload Cryomodules at S10 adit and transport to final location in Housing | White, Dominique A |
| 181088 | 2 | OTHER | Active | Yes | No | OTHER | 09-MAR-2026 00:00 | 28-AUG-2026 00:00 | Gassner, Georg L. | Align Cryomodules once they've been placed on stands in housing | White, Dominique A |
| 181088 | 3 | OTHER | Active | Yes | No | OTHER | 09-MAR-2026 00:00 | 28-AUG-2026 00:00 | Johnson, Scot L. | Support Unload of Cryomodules at S10 adit during LDT | White, Dominique A |
| 181088 | 4 | MFDM | Active | Yes | No | OTHER | 09-MAR-2026 00:00 | - | Chase, Randall | Cryomodule interconnect welds and inspection | White, Dominique A |
| 181088 | 7 | MFDM | Active | Yes | No | OTHER | 01-JUN-2026 00:00 | - | White, Dominique A | Pressure test of cryo lines after interconnect welds are complete | White, Dominique A |
| 181088 | 8 | MFDM | Active | Yes | No | OTHER | 01-JUN-2026 00:00 | - | Atkinson, Richard D. | Cryomodule interconnect close up | White, Dominique A |
| 181347 | 1 | RPFO | Active | Yes | No | OTHER | 30-NOV-2026 00:00 | 18-DEC-2026 00:00 | Najera, Melanie A | Walkthrough to update RP survey maps for LINAC West at end of LDT. | - |
| 181504 | 1 | MFDM | Active | Yes | No | OTHER | 12-JAN-2026 00:00 | - | Atkinson, Richard D. | After demo, Install LCLS-II stands and devices in S10 West. | - |
| 181504 | 2 | MFDV | Active | Yes | No | OTHER | 09-FEB-2026 00:00 | - | Diaz, Gregory L. | Install drift tubes after stands installation | - |
| 181958 | 18 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Cable Demo P08-06 to P08-17 | Hernandez, Joel |
| 181958 | 19 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Cable Demo S9US (Penetrations 1-8) | Hernandez, Joel |
| 181958 | 20 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Cable Demo S9DS Penetrations 9-16 | Hernandez, Joel |
| 181958 | 21 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Cable Demo S10 (tunnel) | Hernandez, Joel |
| 181958 | 22 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Cable Demo S7-PF Tunnel | Hernandez, Joel |
| 181958 | 23 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Cable Demo S7-UHV (Tunnel Demo) | Hernandez, Joel |
| 181958 | 24 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Cable Demo P08-01 to P08-5 (Tunnel) | Hernandez, Joel |
| 181958 | 25 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 13-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray P08-06 to P08-17 | Hernandez, Joel |
| 181958 | 26 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 13-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray S9 Up Stream | Hernandez, Joel |
| 181958 | 27 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray S9 Down Stream | Hernandez, Joel |
| 181958 | 28 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray S10 Tunnel | Hernandez, Joel |
| 181958 | 29 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray S7-PF S7P8-7P10 | Hernandez, Joel |
| 181958 | 30 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 27-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray S7-UHV S7P9-S8-1 | Hernandez, Joel |
| 181958 | 31 | CBL | Active | Yes | No | OTHER | 29-JAN-2026 00:00 | 27-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray P08-01 to P08-05 | Hernandez, Joel |
| 181958 | 32 | CBL | Active | Yes | No | OTHER | 01-FEB-2026 00:00 | 27-FEB-2026 00:00 | Hernandez, Joel | Shift Cable Tray @ PPS Gate between Sector 7-8 tunnel. Contractors need to remove wall before this work can be | Hernandez, Joel |
| 182313 | 1 | RPFO | Active | Yes | No | OTHER | 09-MAR-2026 00:00 | 11-MAR-2026 00:00 | Najera, Melanie A | Replace RP ozone sensor in Sector 2 | Najera, Melanie A |
| 182347 | 1 | MFDM | Active | Yes | No | OTHER | 04-MAY-2026 00:00 | 30-SEP-2026 00:00 | Atkinson, Richard D. | Remove, inspect for proper fit, clean, and reinstall the following Cryomodule CD valves - CM01, CM08, CM10, CM | Galindo, Alondra |
| 178464 | 2 | SSED | Active | Yes | No | PPS | 17-DEC-2025 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install green light at S10 Adit entrance. PPS (outside) | Nicolas, Ludovic Y. |
| 181584 | 1 | CBL | Active | Yes | Yes | PPS | 18-DEC-2025 00:00 | - | Hernandez, Joel | Cut existing cable ID LHE07780 below and above penetration Sector 8 Penetration 12. Install New junction box i | Hernandez, Joel |
| 181584 | 2 | CBL | Active | Yes | Yes | PPS | 17-DEC-2025 00:00 | - | Hernandez, Joel | Cut existing cable ID's LHE07781,2,3 (12PR Trunks) Install new junction box above and below Sector 9 P 4 Insta | Hernandez, Joel |
| 181584 | 4 | CBL | Active | Yes | No | PPS | 18-DEC-2025 00:00 | - | Hernandez, Joel | Install New RP Radiation Monitoring cable Sector 10 P8 | Hernandez, Joel |
| 181584 | 6 | CBL | Active | Yes | No | PPS | 06-JAN-2026 00:00 | - | Hernandez, Joel | Install New Copper Pipe at Sector 10 P8, S7P11, S8P12 | Hernandez, Joel |
| 180000 | 2 | CTL | Active | Yes | No | PROF | 15-JUL-2026 00:00 | - | Curtis, Courtney A. | Reconnect and test OTRDOG in new location | Curtis, Courtney A. |
| 181205 | 10 | CTL | Active | Yes | No | PROF | 05-JAN-2026 00:00 | - | Mock, Jeremy A | Disconnect MPS cables in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181205 | 5 | CTL | Active | Yes | No | PROF | 07-JAN-2026 00:00 | - | Lam, Briant K. | Disconnect Magnets cables in LCLS-II DOGLEG in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181205 | 7 | CTL | Active | Yes | No | PROF | 05-JAN-2026 00:00 | - | Sanchez, Daniel | Disconnect Timing cables in LCLS-II DOGLEG in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181508 | 1 | CTL | Active | Yes | No | PROF | 16-MAR-2026 00:00 | - | Curtis, Courtney A. | TS&R OTRDG02 Target Lamp | Curtis, Courtney A. |
| 181509 | 1 | CTL | Active | Yes | No | PROF | 16-MAR-2026 00:00 | - | Curtis, Courtney A. | TS&R Voltage issue with YAGH2 to filters. Suspect issue with control cable when connected to optics box. | Curtis, Courtney A. |
| 159755 | 2 | AMRF | Active | Yes | No | RF | 13-JUL-2026 00:00 | 22-JUL-2026 00:00 | Nguyen, Binh | Downtime- LCLS-II Gun LLRF install new RF interlock chassis for EPS.
Downtime job. | Benwell, Andy |
| 159755 | 4 | AMRF | Active | No | No | RF | 20-APR-2026 00:00 | 11-DEC-2026 00:00 | Benwell, Andy | Install parallel LLRF system for the LCLS-II gun, test during downtime | Benwell, Andy |
| 170265 | 1 | AMRF | Active | Yes | No | RF | 18-JUN-2026 00:00 | - | Benwell, Andy | Downtime- CM28 Cav 1 - Stepper motor limit switch bypass - if direction of detune is confidently known, we cou | Benwell, Andy |
| 175924 | 4 | AMRF | Active | Yes | No | RF | 12-JAN-2026 00:00 | 15-MAY-2026 00:00 | Kelly, Richard S | CM33 - 8 try to detune cavity approximately 30 kHz with a box at the CM during a PAMM. This will be a 'blind' | Benwell, Andy |
| 177036 | 2 | AMRF | Active | Yes | No | RF | 01-JUN-2026 00:00 | 01-NOV-2026 00:00 | Benwell, Andy | Some SC RF BSA quantities are not showing up in the EPICS control screens. | Benwell, Andy |
| 179591 | 1 | AMRF | Active | Yes | No | RF | 08-JAN-2026 00:00 | 28-JAN-2026 00:00 | Nguyen, Phillip P. | CM09 Cavity 8, monitor stepper and coupler temperatures. Check for any poor connections leading to temperatur | - |
| 181185 | 1 | AMRF | Active | Yes | No | RF | 14-JAN-2026 00:00 | 28-JAN-2026 00:00 | Kelly, Richard S | CM27-3 faulted on STT. Set offline.
Check with the test box. | - |
| 181313 | 1 | AMRF | Active | Yes | No | RF | 16-DEC-2025 00:00 | 16-JAN-2026 00:00 | Kelly, Richard S | CMH2 Cav2 - check SSA characterization for signs of more than enough drive and lower SSA drive max to match.
| - |
| 181494 | 1 | AMRF | Active | Yes | No | RF | 07-JAN-2026 00:00 | 14-JAN-2026 00:00 | Nguyen, Phillip P. | CM09-8 stepper temperature out of tolerance and offline.
Check with the test box.
This station has a hist | - |
| 181525 | 1 | AMRF | Active | Yes | No | RF | 16-DEC-2025 00:00 | - | Nguyen, Phillip P. | CM 34 RFS2B and PRCB check and reseat the RFS2B and PRCB connection at K1. A poor connection here can possibl | Benwell, Andy |
| 167810 | 2 | CTL | Active | Yes | No | SRF | 01-APR-2026 00:00 | 31-JUL-2026 00:00 | Wilson, Andrew | Verify operation as part of calibration after sensor replacement let by Rob Coy. | Wilson, Andrew |
| 177548 | 1 | CTL | Active | No | No | SRF | 02-MAR-2026 00:00 | 29-MAY-2026 00:00 | Wilson, Andrew | Troubleshoot CM29 cavity 5 heater system
-Look for loose wiring on power supply and P24
-Measure voltage of | Wilson, Andrew |
| 177548 | 2 | CTL | Active | No | No | SRF | 02-MAR-2026 00:00 | 29-MAY-2026 00:00 | Wilson, Andrew | Replace CM29 cavity 5 heater power supply
Coordinate with cryo control room. LOTO (cord and plug) power to CM | Wilson, Andrew |
| 177548 | 3 | CTL | Active | No | No | SRF | 02-MAR-2026 00:00 | 29-MAY-2026 00:00 | Wilson, Andrew | Troubleshoot CM26 cavity 3 heater system
-Look for loose wiring on power supply and P23
-Measure voltage of | Wilson, Andrew |
| 177548 | 4 | CTL | Active | No | No | SRF | 02-MAR-2026 00:00 | 29-MAY-2026 00:00 | Wilson, Andrew | Replace CM26 cavity 3 heater power supply
Coordinate with cryo control room. LOTO (cord and plug) power to CM | Wilson, Andrew |
| 177548 | 5 | CTL | Active | Yes | No | SRF | 02-MAR-2026 00:00 | 29-MAY-2026 00:00 | Wilson, Andrew | Troubleshoot CM04 cavity 7 heater system
-Look for loose wiring on power supply and P24
-Measure voltage of | Wilson, Andrew |
| 177548 | 6 | CTL | Active | Yes | No | SRF | 02-MAR-2026 00:00 | 29-MAY-2026 00:00 | Wilson, Andrew | Replace CM04 cavity 7 heater power supply
Coordinate with cryo control room. LOTO (cord and plug) power to CM | Wilson, Andrew |
| 181277 | 1 | MFDV | Active | Yes | No | SRF | 09-FEB-2026 00:00 | 14-FEB-2026 00:00 | Coy, Robert | Remove BLA from CM33-34 interconnect | Coy, Robert |
| 178575 | 1 | MFDV | Active | Yes | No | VACUUM | 16-DEC-2025 00:00 | 06-JAN-2026 00:00 | Diaz, Gregory L. | L1B, He Scav Cart L1B -0295 CDA interlock fault troubleshooting due to either sensor is not functioning proper | Diaz, Gregory L. |
| 178576 | 1 | MFDV | Active | Yes | No | VACUUM | 16-DEC-2025 00:00 | 06-JAN-2026 00:00 | Diaz, Gregory L. | L3, He Scav Carts CDA interlock faults troubleshooting at 6 locations. There was an unexpected function of the | Diaz, Gregory L. |
| 180219 | 2 | MFDM | Active | Yes | No | VACUUM | 05-JAN-2026 00:00 | 31-JAN-2026 00:00 | Atkinson, Richard D. | Support removal of L3 Extension during LDT. | White, Dominique A |
| 181088 | 5 | MFDV | Active | Yes | No | VACUUM | 09-MAR-2026 00:00 | - | Reed, Dentell V. | Leak Check of interconnects after weld | White, Dominique A |
| 181088 | 6 | MFDV | Active | Yes | No | VACUUM | 09-MAR-2026 00:00 | - | Reed, Dentell V. | Particle Free interconnects of Cryomodules | White, Dominique A |
| 182521 | 1 | CTL | Active | Yes | No | VACUUM | 06-FEB-2026 00:00 | - | Karimian, Sherwin | Connect existing vacuum control cables for Pumpcart 1195 in Sector 3 | Karimian, Sherwin |
| 180841 | 17 | CEFOM | Active | Yes | No | WATER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Harrison, William Gerard | Facilities to support the demo of MAG LCW in the following areas -
1. Sector 07 of the tunnel/housing
2. Sec | Perez, Jase |
| 180841 | 30 | CEFFT | Active | No | No | WATER | 09-FEB-0026 00:00 | 13-MAR-0026 00:00 | Perez, Jase | Requesting fire tech support to drain the fire suppression line at penetrations 7-6 and 7-14 that will impact | Perez, Jase |
| 182310 | 1 | CEFOM | Active | Yes | No | WATER | 02-FEB-2026 00:00 | 01-APR-2026 00:00 | Cardenas, Fernando | Demolition / Removal of the now-disconnected and abandoned piping and associated supports in S9.
Piping is | Marshall, Dave |
| 182310 | 2 | CEFOM | Active | Yes | No | WATER | 02-FEB-2026 00:00 | 01-APR-2026 00:00 | Cardenas, Fernando | The sector 5 mechanical alcove supplies Mag LCW water both east and west of the alcove. Now that LPI has modi | Marshall, Dave |
| 182310 | 3 | CEFOM | Active | Yes | No | WATER | 02-FEB-2026 00:00 | 01-APR-2026 00:00 | Cardenas, Fernando | Demolition / Removal of the now-disconnected and abandoned piping and associated supports in S5 to S7.
The | Marshall, Dave |
| 181443 | 1 | CTL | Complete | Yes | No | ATCA | 16-DEC-2025 00:00 | 09-JAN-2026 00:00 | Vu, Minh K | SEC06-SEC10 Install NIC card 10G for all 8100s that connect to ATCA - SC ONLY
1. L2KG06-6044 P4 cpu-l3b-sp | Vu, Thuy Mong |
| 181473 | 1 | CTL | Complete | Yes | No | ATCA | 17-DEC-2025 00:00 | 19-DEC-2025 00:00 | Urbina, Maria | SEC07-SEC10 Power down ATCAs and servers for HE work., except RF HUT in SEC10 (SC only, not FACET) | Vu, Thuy Mong |
| 179953 | 7 | HWPH | Complete | Yes | No | BLEN | 12-JAN-2026 00:00 | 30-JAN-2026 00:00 | Kruchinin, Konstantin | Install a box to house the BLEN amplifiers and a power supply under the beamline, with lead shielding. | Kruchinin, Konstantin |
| 181446 | 1 | PEM | Complete | Yes | No | DC MAGNETS | 16-DEC-2025 00:00 | 09-JAN-2026 00:00 | Herr, Chang | LOTO LCLS-II EXT and DOGLEG magnet power supplies in racks -
L2KG08-51 thru 56
L2KG14-51 thru 52
| Lam, Briant K. |
| 181446 | 2 | PEM | Complete | Yes | Yes | DC MAGNETS | 17-DEC-2025 00:00 | 09-JAN-2026 00:00 | Herr, Chang | Airgap LCLS-II DOGLEG magnets BRB1 and BRB2 only. These are BCS monitored, requiring a RSWCF. | Lam, Briant K. |
| 181446 | 3 | PEM | Complete | Yes | No | DC MAGNETS | 05-JAN-2026 00:00 | 16-JAN-2026 00:00 | Herr, Chang | Airgap LCLS-II EXT magnets from QX01 to BCXDLU4. Orange vest required and procedure required. Full list -
Q | Lam, Briant K. |
| 181446 | 4 | PEM | Complete | Yes | No | DC MAGNETS | 17-DEC-2025 00:00 | 09-JAN-2026 00:00 | Lam, Briant K. | Airgap LCLS-II DOGLEG magnets from YCDOG0 to YCBP12, excluding BRB1 and BRB2 (in job 2 with RSWCF requirements | Lam, Briant K. |
| 181446 | 5 | PEM | Complete | Yes | No | DC MAGNETS | 12-JAN-2026 00:00 | 16-JAN-2026 00:00 | Erebia, Agapito Pete | Valve off and disconnect PCW to racks L2KG08-51 thru 56. Blow out residual water in rack heat exchanger. | Lam, Briant K. |
| 181975 | 1 | CEFE | Complete | Yes | No | ELECTRICITY | 12-JAN-2026 00:00 | - | Hart, Jeff | Provide temporary power to S10 Adit tunnel through penetration 10-8 conduit. | Hart, Jeff |
| 181179 | 2 | CTL | Complete | Yes | No | MOVERS | 06-JAN-2026 00:00 | - | Bong, James | Remove J-panel and pigtails for CEDOG in S07 and store for reinstallation. | Bong, James |
| 181579 | 1 | CBL | Complete | Yes | No | NETWORK INFRASTRUCTURE | 16-DEC-2025 00:00 | - | Hernandez, Joel | Job #HENT11 for Arjun and Charley
Installation of x7 network cables and terminations from KF10-8A to Z=3106. | - |
| 181579 | 2 | CBL | Complete | Yes | No | NETWORK INFRASTRUCTURE | 16-DEC-2025 00:00 | - | Hernandez, Joel | Job #HENT10 Installation of Network cables from Sector 10 Gallery rack LKG10 rack #24 to Sector 10 Adit for Do | - |
| 181581 | 1 | CBL | Complete | Yes | No | NETWORK INFRASTRUCTURE | 16-DEC-2025 00:00 | - | Hernandez, Joel | Installation of Network Cables from Sector 10 gallery and Alcove to Sector 10 Adit. These cables are being put | Hernandez, Joel |
| 180841 | 18 | SSED | Complete | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove unused ODM system panels inside L2KA08-05 PPS Rack (back). | Perez, Jase |
| 181577 | 1 | SSED | Complete | Yes | Yes | ODM | 09-JAN-2026 00:00 | - | Nicolas, Ludovic Y. | Power down the Linac West ODM prior to the S7-S10 rack power outage. | Cyterski, Matt |
| 180219 | 3 | MFDV | Complete | No | No | OTHER | 16-DEC-2025 00:00 | 23-JAN-2026 00:00 | Diaz, Gregory L. | Vent beam-line from gate valve at beginning of dogleg to gate valve at beginning of S12. | White, Dominique A |
| 180219 | 5 | MFDM | Complete | Yes | Yes | OTHER | 16-DEC-2025 00:00 | 09-JAN-2026 00:00 | Atkinson, Richard D. | S10 wall shielding installation.
Once the UHV beam-line is vented, vacuum team will remove one beam-line se | White, Dominique A |
| 180658 | 1 | CBL | Complete | Yes | No | OTHER | 16-DEC-2025 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Stage strut and hardware in Sector 7-10, for cable shifting work as per HE drawings. | Hernandez, Joel |
| 181116 | 1 | RPFO | Complete | Yes | No | OTHER | 16-DEC-2025 00:00 | 19-DEC-2025 00:00 | Najera, Melanie A | Remove ozone sensor box from tunnel at 07-11 | Najera, Melanie A |
| 181117 | 1 | RPFO | Complete | Yes | Yes | OTHER | 16-DEC-2025 00:00 | 19-DEC-2025 00:00 | Harris, Anne M. | Remove RDM at Sec 7 (before SEC 7-10 demo) with it's pre-amp from the housing at Sec 7-11, and remove the disp | Harris, Anne M. |
| 181242 | 1 | RPFO | Complete | Yes | No | OTHER | 16-DEC-2025 00:00 | 19-DEC-2025 00:00 | Harris, Anne M. | Fiber optic cables (in bags with thermometer) for radiation hardness testing at SEC 1, across fromPC0 shieldin | Harris, Anne M. |
| 181375 | 1 | MFDM | Complete | Yes | Yes | OTHER | 09-JAN-2026 00:00 | 17-JAN-2026 00:00 | Atkinson, Richard D. | Remove shield caps from penetrations listed below and coordinate with riggers to move the caps to B15 parking | Loo, Alysson |
| 181432 | 1 | OTHER | Complete | Yes | No | OTHER | 16-DEC-2025 00:00 | 16-DEC-2025 00:00 | Griffie, Larry | AD118P on 12/16 @ 1pm | Griffie, Larry |
| 181432 | 2 | OTHER | Complete | Yes | No | OTHER | 17-DEC-2025 00:00 | 17-DEC-2025 00:00 | Griffie, Larry | AD118P on 12/17 @ 9am | Griffie, Larry |
| 181958 | 1 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 23-JAN-2026 00:00 | Hernandez, Joel | Gallery Controls Demo P08-06 to P08-17. All cables | Hernandez, Joel |
| 181958 | 2 | CBL | Complete | Yes | No | OTHER | 14-JAN-2026 00:00 | 23-JAN-2026 00:00 | Hernandez, Joel | Tray Demo Sec 8 P6 | Hernandez, Joel |
| 181958 | 3 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 23-JAN-2026 00:00 | Hernandez, Joel | Rack Relocation Sec 8 P6, racks will go back to B15 to be reconfigured | Hernandez, Joel |
| 181958 | 4 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Tray Demo Sec 8 P12 | Hernandez, Joel |
| 181958 | 5 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Rack Relocation Sec 8 P12 | Hernandez, Joel |
| 181958 | 6 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Gallery Controls Demo S9US (Pen Sec 9 P1-8) All system cables from gallery racks and alcoves to penetrations | Hernandez, Joel |
| 182486 | 1 | AEG | Complete | No | No | OTHER | 05-FEB-2026 00:00 | 05-FEB-2026 00:00 | Gassner, Georg L. | Layout the location of the new PPS gate at S7 | Gassner, Georg L. |
| 180219 | 6 | MFDM | Complete | Yes | Yes | PPS | 17-DEC-2025 00:00 | 07-JAN-2026 00:00 | Diaz, Gregory L. | Sec 10 adit concrete double doors - remove the floor threshold shielding shims. This is needed to remove the f | White, Dominique A |
| 181118 | 1 | RPFO | Complete | Yes | Yes | PPS | 07-JAN-2026 00:00 | 09-JAN-2026 00:00 | Harris, Anne M. | Remove neutron and photon BSOICS (BSOLNW10 and BSOLNW11n) located in Sec 7 (prior to Sec 7- Sec10 demo). | Harris, Anne M. |
| 181584 | 3 | CBL | Complete | Yes | No | PPS | 16-DEC-2025 00:00 | - | Hernandez, Joel | Remove Copper Pipe at penetration's Sec 7P11 & Sec 8 P12 | Hernandez, Joel |
| 181584 | 5 | CBL | Complete | Yes | No | PPS | 17-DEC-2025 00:00 | - | Hernandez, Joel | Remove & Save RP Cable at Sector 7 P12 Near RDM (Grey, Do Not cut) | Hernandez, Joel |
| 181205 | 1 | CTL | Complete | Yes | No | PROF | 16-DEC-2025 00:00 | - | Curtis, Courtney A. | Disconnect Profile Monitor cables in LCLS-II DOGLEG area support HE Demo for S7-10 | Saraf, Shweta |
| 181205 | 2 | CTL | Complete | Yes | No | PROF | 17-DEC-2025 00:00 | 19-DEC-2025 00:00 | Murphy, James Matthew | Disconnect BPM cables in LCLS-II DOGLEG area in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181205 | 3 | CTL | Complete | Yes | No | PROF | 05-JAN-2026 00:00 | - | Sanchez, Daniel | Disconnect Vacuum cables in S7-10 in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181205 | 4 | CTL | Complete | Yes | No | PROF | 05-JAN-2026 00:00 | - | Bong, James | Disconnect Collimator (CEDOG) and fast wire scanner (WSBP1) long haul cables in LCLS-II DOGLEG in support of S | Saraf, Shweta |
| 181205 | 6 | CTL | Complete | Yes | No | PROF | 05-JAN-2026 00:00 | - | Karimian, Sherwin | Disconnect Temperature cables in S7-10 in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181205 | 8 | CTL | Complete | Yes | No | PROF | 05-JAN-2026 00:00 | - | Shetty, Arjun | Disconnect Network cables in support of S7-10 HE Demo Work | Saraf, Shweta |
| 181205 | 9 | CTL | Complete | Yes | No | PROF | 18-DEC-2025 00:00 | - | Taylor, Zack | Disconnect and mark the following cables in the LLRF racks in support of S7-10 HE Demo Work.
CM31-CAVITY 1-4 | Saraf, Shweta |
| 181585 | 1 | CBL | Complete | Yes | No | RADIOS | 17-DEC-2025 00:00 | 27-FEB-2026 00:00 | Hernandez, Joel | Relocate Antenna Amplifier from Gallery to Tunnel in Sector 9 P12 Install New cables to existing long hauls, m | - |
| 144826 | 91 | CTL | Complete | No | No | VACUUM | 16-DEC-2025 00:00 | - | Karimian, Sherwin | Replace TCP 350 Turbo Controller in L2KA03-0130 due to Error 993 | Karimian, Sherwin |
| 180219 | 1 | MFDV | Complete | Yes | No | VACUUM | 05-JAN-2026 00:00 | 31-JAN-2026 00:00 | Diaz, Gregory L. | Removal of Particle free beamline (L3 Extension) from housing during LDT. Only valve extension will remain in | White, Dominique A |
| 181460 | 1 | CTL | Complete | Yes | No | VACUUM | 16-DEC-2025 00:00 | - | Karimian, Sherwin | Zero out Pirani channels on gauge controllers to resolve readback issue | Karimian, Sherwin |
| 181561 | 1 | CBL | Complete | Yes | No | VACUUM | 16-DEC-2025 00:00 | - | Hernandez, Joel | Install temporary CAT6 cable from South wall, at Z location near Pen 7-06, to North wall near L3 endcap device | Karimian, Sherwin |
| 181561 | 2 | CTL | Complete | Yes | No | VACUUM | 16-DEC-2025 00:00 | - | Karimian, Sherwin | Temporary setup of DIGI, ion pump controller, and gauge controller in housing to support local monitoring of L | Karimian, Sherwin |
| 181561 | 3 | CTL | Complete | Yes | No | VACUUM | 16-DEC-2025 00:00 | - | Karimian, Sherwin | Lock close the following PF vacuum valves to support CM warm-up and LDT work -
VVPG -GUNB -941
VVPG -HTR - | Karimian, Sherwin |
| 181473 | 2 | CTL | Drop | Yes | No | ATCA | 08-JAN-2026 00:00 | 09-JAN-2026 00:00 | Tang, Van | SEC07-SEC10 Power restart ATCAs and servers after HE work, except RF HUT in SEC10 (SC only, not FACET) | Vu, Thuy Mong |
| 181091 | 1 | ME | Drop | Yes | No | DC MAGNETS | 16-DEC-2025 00:00 | 17-DEC-2025 00:00 | Anderson, Scott D. | Measure the water flow for DOG magnets that are not planned to go to the magnetic measurements shop during the | Anderson, Scott D. |
| 182223 | 1 | ADSO | Drop | No | Yes | OTHER | 22-JAN-2026 00:00 | 23-JAN-2026 00:00 | Andrews, Anthony | Unlock ADSO lock at shed above the S10 Adit. Access is needed for contractors. | Andrews, Anthony |
| 180000 | 1 | CTL | Drop | No | No | PROF | 16-DEC-2025 00:00 | 30-JAN-2026 00:00 | Curtis, Courtney A. | Test OTRDOG status, then disconnect cables from device and in rack for device and rack move for HE---Dropped f | Curtis, Courtney A. |
| 178043 | 1 | PEM | Roll | Yes | No | DC MAGNETS | 21-JUL-2025 00:00 | 29-AUG-2025 00:00 | Lam, Briant K. | CM22 - Inspect magnet power and voltage tap wiring to troubleshoot false quench faults. | Lam, Briant K. |
| 178043 | 2 | PEM | Roll | Yes | No | DC MAGNETS | 21-JUL-2025 00:00 | 22-AUG-2025 00:00 | Lam, Briant K. | CM30 - Inspect magnet power and voltage tap wiring to troubleshoot false quench faults. | Lam, Briant K. |
LINAC WEST GALLERY |
| Shop |
| 180841 | 4 | CEFOM | Active | Yes | No | COMPRESSED AIR | 19-JAN-2026 00:00 | 06-MAR-2026 00:00 | Perez, Jase | Demo of the existing CDA filter at penetration 7-14 and provide LOTO for compressed air in Sector 07 and 08 of | Perez, Jase |
| 181114 | 2 | MFDM | Active | Yes | Yes | CRYO DISTRIBUTION | 23-MAR-2026 00:00 | 31-MAR-2026 00:00 | Ki, Taekyung | VTL (Vertical Transfer Line) 3 Installation | Hiatt, Tommy |
| 180829 | 1 | CEFE | Active | Yes | No | ELECTRICITY | 12-JAN-2026 00:00 | 12-JAN-2026 00:00 | Hart, Jeff | HE LDT - LOTO support for EED Sector 7-10 rack & cable demo
Request is for F&O electricians develop an ener | Hart, Jeff |
| 180841 | 14 | SSED | Active | Yes | No | ODM | 12-JAN-2026 00:00 | 16-JAN-2026 00:00 | Nicolas, Ludovic Y. | Inspect, tag and disconnect S07 ODM cables that will need to be removed and/or reused. | Perez, Jase |
| 180841 | 16 | SSED | Active | Yes | No | ODM | 12-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Inspect ODM racks wiring in the alcove of S01-S08 to update drawings and serve a basis for the design of HE OD | Perez, Jase |
| 180841 | 1 | OTHER | Active | Yes | Yes | ODM | 02-FEB-2026 00:00 | 27-MAR-2026 00:00 | Perez, Jase | Sector 07 AHU will get demo'd during the long downtime.
Ductwork will need to get demo'd and existing circ | Perez, Jase |
| 181557 | 2 | SSED | Active | Yes | No | ODM | 13-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install UltraFlow Duo Oxigraf in S00. | Nicolas, Ludovic Y. |
| 181557 | 3 | SSED | Active | Yes | No | ODM | 13-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove S02 Original Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 4 | SSED | Active | Yes | No | ODM | 13-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Install S02 UltraFlow Duo Oxigraf | Nicolas, Ludovic Y. |
| 181557 | 5 | SSED | Active | Yes | No | ODM | 15-JAN-2026 00:00 | 15-DEC-2026 00:00 | Nicolas, Ludovic Y. | Remove S03 Original Oxigraf | Nicolas, Ludovic Y. |
| 181593 | 1 | MFDM | Active | Yes | Yes | ODM | 19-DEC-2025 00:00 | 05-OCT-2026 00:00 | Atkinson, Richard D. | Open louvers in ODH ducting at S1P18, S2P18, and S4P18. Louvers will be returned to closed positions after the | Atkinson, Richard D. |
| 180459 | 4 | OTHER | Active | Yes | No | OTHER | 22-JAN-2026 00:00 | 30-APR-2026 00:00 | Checchin, Mattia | Install SSA and waveguide frames anchors in S7 through S9.
SOW steps -
- drill and install anchors (contract | Checchin, Mattia |
| 180459 | 5 | OTHER | Active | Yes | No | OTHER | 05-JAN-2026 00:00 | 23-JAN-2026 00:00 | Checchin, Mattia | Install penetration shield bases anchors in P9-05 and P7-17 -
- drill and install anchors (contractor)
- spe | Checchin, Mattia |
| 180459 | 6 | OTHER | Active | Yes | Yes | OTHER | 02-FEB-2026 00:00 | 30-APR-2026 00:00 | Checchin, Mattia | Install anchors and hangers for waveguide on S7 housing N-wall
- install anchors (contractor)
- special insp | Checchin, Mattia |
| 180841 | 5 | CEFOM | Active | Yes | No | OTHER | 02-FEB-2026 00:00 | 27-FEB-2026 00:00 | Perez, Jase | Subcontractor to install new He Guard for future cryomodules
FWKD0114957 | Perez, Jase |
| 181375 | 2 | OTHER | Active | Yes | Yes | OTHER | 28-JAN-2026 00:00 | 30-JAN-2026 00:00 | Loo, Alysson | First set of HPRF Penetration Installations at P08-05 and P09-06. Work is part of RSWCF 10672 | Loo, Alysson |
| 181375 | 3 | OTHER | Active | Yes | Yes | OTHER | 24-FEB-2026 00:00 | 03-MAR-2026 00:00 | Loo, Alysson | Second set of HPRF penetration installations -
P07-06
P07-07
P07-09
P07-10
P07-11
P07-12
P07-13
P07-14 | Loo, Alysson |
| 181958 | 12 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 13-FEB-2026 00:00 | Hernandez, Joel | Sector 10 Gallery Controls Demo | Hernandez, Joel |
| 181958 | 13 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 13-FEB-2026 00:00 | Hernandez, Joel | Gallery Controls Demo S7-PF (Down Stream) (Penetrations 8-16) | Hernandez, Joel |
| 181958 | 14 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Gallery Controls Demo S7-UHV (Up Stream) (Penetrations 1-15) | Hernandez, Joel |
| 181958 | 15 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 13-FEB-2026 00:00 | Hernandez, Joel | Gallery Controls Demo P08-01 to P08-05 | Hernandez, Joel |
| 181958 | 16 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 13-FEB-2026 00:00 | Hernandez, Joel | Tray Demo Sec 8 P4 Gallery | Hernandez, Joel |
| 181958 | 17 | CBL | Active | Yes | No | OTHER | 20-JAN-2026 00:00 | 20-FEB-2026 00:00 | Hernandez, Joel | Rack Relocation Sec 8 P4 (Gallery) | Hernandez, Joel |
| 180477 | 4 | CTL | Active | Yes | No | RADIOS | 15-JAN-2026 00:00 | - | Bong, James | Test tunnel radio amplifiers in S01 Alcove and gallery, S03 & S06 Gallery, and S09 Tunnel. | Bong, James |
| 180477 | 5 | CTL | Active | Yes | No | RADIOS | 16-JAN-2026 00:00 | - | Bong, James | Replace failing 24VDC Power Supply in S06 Tunnel Radio Amplifier Chassis. | Bong, James |
| 180871 | 1 | AMRF | Active | Yes | No | RF | 12-JAN-2026 00:00 | 16-JAN-2026 00:00 | Kelly, Richard S | H2-3 still has a large mismatch between forward and reverse power after PAMM work.
Replace the directional | Kelly, Richard S |
| 181255 | 1 | AMRF | Active | Yes | No | RF | 16-DEC-2025 00:00 | 16-JAN-2026 00:00 | Kelly, Richard S | CM4A, 14B, 22B, and 34B NRP pressure.
Adjust all NERP pressure gauges to 2 PSI. | - |
| 180841 | 10 | CEFFT | Active | Yes | No | WATER | 12-JAN-2026 00:00 | 28-FEB-2026 00:00 | Perez, Jase | Fire suppression lines in penetrations 7-06 and 7-14 will require relocation.
A drain down of the system a | Perez, Jase |
| 181722 | 1 | CTL | Complete | Yes | No | ATCA | 17-DEC-2025 00:00 | 30-JAN-2026 00:00 | Vu, Minh K | L2KG01-20 P23 cpu-col0-sp01 (M308-8100) is not responding to ping and there is no response on the console. Inv | Vu, Thuy Mong |
| 180841 | 2 | CEFOE | Complete | Yes | No | ELECTRICITY | 16-DEC-2025 00:00 | 18-DEC-2025 00:00 | Perez, Jase | Sector 10 - MCC4P10G will require a LOTO for the tie in of a new circuit.
New circuit will serve the new pa | Perez, Jase |
| 181179 | 1 | CTL | Complete | Yes | No | MOVERS | 06-JAN-2026 00:00 | - | Bong, James | Remove Aerotech Controller, LVDT Conditioner, Pigtail cables, and termination panel from rack L2KA07-05 | Bong, James |
| 180841 | 11 | SSED | Complete | Yes | No | ODM | 05-JAN-2026 00:00 | 09-JAN-2026 00:00 | Nicolas, Ludovic Y. | Power down Linac ODM system S00-S07 for LDT. | Perez, Jase |
| 180841 | 15 | SSED | Complete | Yes | No | ODM | 12-JAN-2026 00:00 | 16-JAN-2026 00:00 | Nicolas, Ludovic Y. | Dismount and remove S07 oxigraf and its cabinet from Penetration 07-11. | Perez, Jase |
| 181114 | 1 | OTHER | Complete | Yes | Yes | OTHER | 17-DEC-2025 00:00 | 18-DEC-2025 00:00 | Loo, Alysson | LCLS-II-HE - Removing temporary shielding blocks for VTL3 in preparation of VTL3 and permanent shielding insta | Hiatt, Tommy |
| 181115 | 1 | RPFO | Complete | Yes | No | OTHER | 16-DEC-2025 00:00 | 19-DEC-2025 00:00 | Najera, Melanie A | Remove AMS and ozone devices inside AMS enclosure at 07-11 in KG (before demo starts) | - |
| 181958 | 10 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Tray Demo Sec 9 P14 | Hernandez, Joel |
| 181958 | 7 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Tray Demo Sec 9 P4 | Hernandez, Joel |
| 181958 | 9 | CBL | Complete | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Gallery Controls Demo S9DS (Pen Sec 9 P9-16) All Systems cables in the gallery | Hernandez, Joel |
| 180841 | 7 | CEFOM | Complete | Yes | No | WATER | 16-DEC-2025 00:00 | 09-JAN-2026 00:00 | Harrison, William Gerard | WO - FWKD0115567
F&O Mechanics to drain down the 30C LCW bypass line when LCLS-SC beamline is off.
| Perez, Jase |
| 180841 | 8 | CEFOM | Complete | Yes | No | WATER | 18-DEC-2025 00:00 | 09-JAN-2026 00:00 | Harrison, William Gerard | WO FWKD0108977
F&O Mechanics to drain down the existing MAG LCW line when LCLS-SC beamline is off.
| Perez, Jase |
| 181958 | 11 | CBL | Drop | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Rack Relocation Sec 9 P14 | Hernandez, Joel |
| 181958 | 8 | CBL | Drop | Yes | No | OTHER | 13-JAN-2026 00:00 | 30-JAN-2026 00:00 | Hernandez, Joel | Rack Relocation Sec 9 P4 | Hernandez, Joel |
| 181061 | 1 | AMRF | Drop | Yes | No | RF | 06-JAN-2026 00:00 | 20-JAN-2026 00:00 | Wack, Bob | CM34-1 is PZt faulted after recovering machine from cryoplant trip.
Replace the piezo PCB. | Wack, Bob |
LINAC MIDDLE GALLERY |
| Shop |
| 180312 | 2 | PEM | Active | Yes | No | KLYSTRON / MODULATOR | 24-JAN-2026 00:00 | 28-FEB-2026 00:00 | Reagan, Jeff | 17-7 T/S and repair Circuit of DC HV meter and Circuit of DQ meter | Chen, Xupeng |
| 181866 | 3 | CBL | Active | Yes | No | OTHER | 09-JAN-2026 00:00 | 29-MAY-2026 00:00 | Mortenson, Paul R. | Sector 11 Penetrations 11-04 and 11-06 remove the Gas Cylinders, Regulator and any support related to it as we | Hugyik, John |
| 179596 | 1 | CEFOE | Complete | Yes | No | ELECTRICITY | 07-OCT-2025 00:00 | 09-OCT-2025 00:00 | Hart, Jeff | The LOTO will affect LCLS II racks L2KG11-4C & 4D
Inspect and repair damaged conduits in sector 11
See Ross | Hart, Jeff |
| 181767 | 1 | AMRF | Complete | No | No | RF | 06-JAN-2026 00:00 | 23-JAN-2026 00:00 | Delgadillo, Eduardo D. | 17-5 - SLED motor not at limit in the tune mode.
Check the operation with the SLED test box. | Kelly, Richard S |
| 181851 | 3 | MFDV | Complete | Yes | No | VACUUM | 02-FEB-2026 00:00 | - | Guerra, Oscar | 14-8 WG valve limit switch replacement. Once completed coordinate with CTL (Alverson) to ensure limit switches | Diaz, Gregory L. |
SOFTWARE |
| Shop |
| 181818 | 1 | SWE | Active | No | No | ALARMS | 18-NOV-2026 00:00 | - | Summers, Tasha | Unbypass the things in the task list if they're ready.
| Summers, Tasha |
|
|
|